Circular No. 6588

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 970228 P. J. Groot and T. J. Galama, University of Amsterdam (UoA); J.

van Paradijs, UoA and University of Alabama, Huntsville; J. Melnick, G. van der Steene, and M. Bremer, European Southern Observatory (ESO); N. Tanvir and J. Bloom, Institute of Astronomy, Cambridge; R. Strom, Netherlands Foundation for Research in Astronomy; J. Telting and R. G. M. Rutten, Isaac Newton Group of Telescopes, La Palma; C. Kouveliotou, Universities Space Research Association; J. in ’t Zand and J. Heise, SRON, Utrecht; E. Costa, M. Feroci, and L. Piro, Istituto di Astrofisica Spaziale, CNR; F. Frontera and G. Zavattini, Universitá di Ferrara; and L. Nicastro and E. Palazzi, Istituto TESRE, CNR, report: “An R-band image (3600-s exposure) made with the ESO New Technology Telescope on Mar. 13.0 UT shows the presence of an extended object at R = 23.8 ± 0.2 within 0′′.2 of the position of the optical transient reported on IAUC 6584. The object is elongated in the north-south direction and measures 1′′.3 × 1′′.0 (FWHM). The point-spread function for the image has 1′′.0 FWHM. This indicates that the object is a galaxy. The brightness of the nearby star (the position end figures for which should have been given on IAUC 6584 as 46^{s}.43 and 53′′.0) is R = 21.5. R- and B-band images made with the Isaac Newton Telescope on Mar. 9.9 show an object at the position of the optical transient whose magnitudes, R = 24.0 and B = 25.4, are likely dominated by the above galaxy. Low-resolution spectroscopy of the nearby star with the ESO 3.6-m telescope shows the presence of TiO features, which together with the color index (V − I = +2.6) indicates that it is an early M-type dwarf unrelated to the optical transient.”

M. R. Metzger, S. R. Kulkarni, S. G. Djorgovski, R. Gal, and C. C. Steidel, California Institute of Technology; and D. A. Frail, National Radio Astronomy Observatory, communicate: “Examination of an R-band image taken with the Keck II 10-m telescope on Mar. 6.32 UT (cf. IAUC 6582) reveals a faint source near the position of the optical variable reported on IAUC 6584. This source appears extended, with R = 24.0 and its center at α = 5^{h}01^{m}46^{s}.59, δ = +11^{o}46′53′′.5 (equinox 2000.0); it is 2′′.7 away from a brighter source with R = 22.4 at position end figures 46^{s}.40, 53′′.3 (± 0′′.2), evidently corresponding to the steady source of Groot et al. (see above). No other sources are detected within 5′′ to R = 25.0. The extended source also appears on an I-band image taken on Mar. 6.19 with the Palomar 1.5m reflector, near the detection limit of I = 21.5. From an R-band image taken on Mar. 11.18 at the Palomar 5-m reflector, we find that the extended source has not varied by more than 0.3 mag in R over the interval.”

1997 March 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green